OpenStudy (math2vcap):

A culture started with 2,000 bacteria. After 4 hours, it grew to 2,600 bacteria. predict how many bacteria will be present after 17 hours. Round you answer to the nearest whole number. P=Ae^kt

OpenStudy (math2vcap):

note: when solving for k round to 4 decimal places

OpenStudy (math2vcap):

\[P=Ae ^{kt}\]

OpenStudy (mariamikayla):

In 4 hours the culture grow up with 2,600-2,000=600 bacteria We use the 3 simple rule: 4hours........600 bacteria 17hours.........x x= 17*600/4 x=17*150 x=2,550 bacteria grew in 17 hours The culture started with 2,000 bacteria so we have to add the ones formed in 17 hours 2,000+2,550= 4,550 bacteria were in total after 17 hours rounded it would be 5,000

OpenStudy (math2vcap):

thanks

OpenStudy (mariamikayla):

You're welcome :D

OpenStudy (triciaal):